OpenStudy (anonymous):

find an equation for the line with the given properties perpendicular to the line x-8y=6 containing the point (2,3) I changed the equation to y=(x+6)/-8 but I'm kind of stuck as to what I'm supposed to next

hero (hero):

How about \[y = -\frac{x}{8} - \frac{3}{4}\]

OpenStudy (anonymous):

so would the slope be -1? where the reciprocal, if it were perpendicular would be 1?

hero (hero):

Remember \[\frac{a + b}{c} = \frac{a}{c} + \frac{b} {c}\]

hero (hero):

The slope is -1/8. Do you see that?

hero (hero):

Maybe if I wrote it this way: \[y = -\frac{1}{8}x - \frac{3}{4}\]

OpenStudy (anonymous):

oooooo.. rioght.. so then the reciprocal would be 8/1??

hero (hero):

Precisely

OpenStudy (anonymous):

so how would that work, to find the "points".. put the equation as y=8x-3?

hero (hero):

You cannot assume what b is. All you know is the slope and the given point: so y = 8x + b insert the given point to find b

OpenStudy (anonymous):

so b=-13?

hero (hero):

Would you like to show how you got b = -13?

OpenStudy (anonymous):

if the points were (2,3) and then equation I had was y=8x+b 3=8(2)+b, solve and then the new equation is y=8x-13?

OpenStudy (anonymous):

or no....?

hero (hero):

Good job

OpenStudy (anonymous):

yay :) perpendicular equations always confused me, and now they've come back to kick my in the butt

hero (hero):

Nah, they're easy. Plus, when you start helping others with similar problems, they get even easier.

OpenStudy (anonymous):

thanks :)

hero (hero):

That's it? No more questions?

OpenStudy (anonymous):

hah, hold on